OpenStudy (anonymous):

when you hold the nozzle of the water horizontally at height of 0.60 meters above the ground, the stream of water lands 4.2 meters from you. If you now aim the nozzle straight up, how high will the stream of water rise ?

OpenStudy (joannablackwelder):

The force that pulls the water to the ground is gravity. So, use the vertical kinematics equations to find the time it takes to reach the ground.

OpenStudy (joannablackwelder):

Then use that time in the horizontal direction to find the original velocity in the x-direction.

OpenStudy (joannablackwelder):

Since that is the horizontal velocity when you hold the hose horizontally, it would also be the vertical velocity if you hold the hose vertically.

OpenStudy (joannablackwelder):

Well, it would give you the initial vertical velocity at least. :)

OpenStudy (anonymous):

i dnt understand actually.. can you tell me with the formula ? please

OpenStudy (joannablackwelder):

Do you understand that the question requires finding the initial vertical velocity of the water to find the height the water rises?

OpenStudy (joannablackwelder):

The final vertical velocity would be 0 since the water would start falling back down from there and the acceleration would be the acceleration of gravity.

OpenStudy (anonymous):

yeah i understand but i am not sure which fornula to apply

OpenStudy (joannablackwelder):

I would use ∆x=vi∆t+1/2a(∆t)^2 to start.

OpenStudy (joannablackwelder):

In the vertical direction. What do you get for the initial velocity in the vertical direction?

OpenStudy (joannablackwelder):

But you will need to use multiple formulas.

OpenStudy (rajat97):

@joannablackwelder is right we need to use the formula ∆x=vi∆t+1/2a(∆t)^2 to start with the y- direction here the initial velocity in the y direction is zero nd the displacement in the y direction is -0.6 m and the accelereation in the y direction is -g(acceleration due to gravity) so we can get the time by putting these values in the equation so you'll get the final equation as \[t = \sqrt{2h/g}\] and here h=0.6 and g=9.8m/s^2 now that you have the time taken for the motion, you can now look at the motion in x-drection the motion in the x-direction is uniform motion as there is no force acting on the water in the x direction so the water will travel with uniform speed in the x direction so we can use the formula distance = speed x time we know the time(\[\sqrt{2h/g}\]) and we know the distance 4.2 m so we can find the initial velocity of the water now that we have the initial velocity of the water, we can now point the nozzel in the upward direction and find the height reached by the water by the kinematical equation v^2 - u^2= 2as here v^2= 0 as the velocity on the topmost point is zero as the direction of motion gets changed at the topmost point. and a=-g so you can solve for s once you've got the value of s , just add 0.6 in your answer if the question says to find the height from the ground by doing all this , i've got the answer 8.38m=7.78m+0.6m i hope this helps you:)

OpenStudy (anonymous):

thanks for the effort but it's wrong

OpenStudy (anonymous):

this is hoemwork question rom chapter motion in one dimension. we are studying physics one

OpenStudy (anonymous):

motion in two dimension i meant

OpenStudy (roadjester):

Okay I'm going to need time to think about that then...

OpenStudy (anonymous):

ok please tell me if u figure out

OpenStudy (carmuz):

As @joannablackwelder said, you have to find Vo, to know the velocity the water comes out of the nozzle.  Ok. First, let's get the time the stream needs to get to the ground: \[h = \frac{ 1 }{ 2 } g t^2 \]\[0.6 = \frac{ 1 }{ 2 } (9.8) t^2\]Here you have: t = 0.35 s The streams needs to go 4.2m in 0.35s (this is very easy -> x=vt), so: Vo = 13.7 m/s You have Vo. Now, all you need is to calculate how high the water will rise : \[V_0 = \sqrt{2gh}\] h = 9.6 m Hope it helps.